OpenStudy (anonymous):

a side of the hexagon and a side of the octagon? I got x= 105

jimthompson5910 (jim_thompson5910):

105 is incorrect

jimthompson5910 (jim_thompson5910):

oh wait, I made a typo

jimthompson5910 (jim_thompson5910):

yeah it's correct

jimthompson5910 (jim_thompson5910):

the interior angle of the hexagon is 180(6-2)/6 = 120 the interior angle of the octagon is 180(8-2)/8 = 135 so, 120+135+x = 360 x = 105 I skipped a few steps, but I'm assuming this is what you have (somewhat)

jimthompson5910 (jim_thompson5910):

yep 360/6 = 60